# ADePT: Adaptive Decomposed Prompt Tuning for Parameter-Efficient Fine-tuning

### v1

@@ Abstract
Prompt Tuning (PT) enables the adaptation of Pre-trained Large Language Models (PLMs) to downstream tasks by optimizing a small amount of soft virtual tokens, which are prepended to the input token embeddings. Recently, Decomposed Prompt Tuning (DePT) has demonstrated superior adaptation capabilities by decomposing the soft prompt into a shorter soft prompt and a pair of low-rank matrices. The product of the pair of low-rank matrices is added to the input token embeddings to offset them. Additionally, DePT achieves faster inference compared to PT due to the shorter soft prompt. However, in this paper, we find that the position-based token embedding offsets of DePT restricts its ability to generalize across diverse model inputs, and that the shared embedding offsets across many token embeddings result in sub-optimization. To tackle these issues, we introduce Adaptive Decomposed Prompt Tuning (ADePT), which is composed of a short soft prompt and a shallow token-shared feed-forward neural network. ADePT utilizes the token-shared feed-forward neural network to learn the embedding offsets for each token, enabling adaptive embedding offsets that vary according to the model input and better optimization of token embedding offsets. This enables ADePT to achieve superior adaptation performance without requiring more inference time or additional trainable parameters compared to vanilla PT and its variants. In comprehensive experiments across 23 natural language processing (NLP) tasks and 4 typical PLMs of different scales, we show that ADePT consistently surpasses the leading parameter-efficient fine-tuning (PEFT) methods, and even outperforms the full fine-tuning baseline in certain scenarios. Code is available at https://github.com/HungerPWAY/ADePT.
@@ 1 Introduction
DePT also learning a short soft prompt $\bm{P}_{s}\in\mathbb{R}^{l_{s}\times d}$ with length $l_{s}$. The input token embeddings $\bm{E}$, after DePT processing, are formulated as $[\bm{P}_{s},\bm{E}+\bm{A}\bm{B}]$. In this paper, we focus on improving the fourth path by providing better offsets for the input token embeddings. Additionally, our improvement of the fourth path is orthogonal to the first three paths and can be integrated with the methods of the first three paths.
@@
  • Extensive evaluations across 23 NLP tasks and 4 PLMs of different scales show that ADePT surpasses leading PEFT methods, including full fine-tuning in certain scenarios.
@@ 2 Related Works
Parameter-Efficient Fine-tuning: PEFT methods can adapt the PLMs to the new target downstream tasks by optimizing only a small amount of parameters, significantly reducing the demand for computational resources. The PEFT methods can be categorized into four types:(1) Adapter and its variants (Houlsby et al., 2019; He et al., 2022; Rücklé et al., 2021; Ivison & Peters, 2022); (2) Low-rank Adaptation (LoRA) and its variants (Hu et al., 2021; Liu et al., 2022; Kopiczko et al., 2024); (3) Prompt Tuning (PT) and its variants (Lester et al., 2021; Li & Liang, 2021; Vu et al., 2022; Asai et al., 2022; Wang et al., 2023; Ma et al., 2022; Xiao et al., 2023; Razdaibiedina et al., 2023; Shi & Lipani, 2024); (4) other methods (Ben Zaken et al., 2022; Guo et al., 2021). The adapter inserts new modules into the transformer blocks (Houlsby et al., 2019). Hyperformer (He et al., 2022) uses a shared hypernetwork to generate task-conditioned adapters, reducing the trainable parameters in multi-task scenarios. AdapterDrop (Rücklé et al., 2021) removes adapters from lower transformer layers during training and inference, which reduces the computation overhead. Hyperdecoders (Ivison & Peters, 2022) generate input-specific adapters using a shared decoder for multi-task scenarios. LoRA (Houlsby et al., 2019) adopts the matrix product of a pair of low-rank matrices to approximate the updates of corresponding parameters. $(\text{IA})^{3}$ (Liu et al., 2022) rescales internal activations only using learned vectors injected into the attention and feedforward modules. Prompt tuning (PT) (Lester et al., 2021)adapts PLMs to the new downstream tasks by optimizing learnable virtual tokens prepared for the model input. LST (Sung et al., 2022) reduces memory requirements by training a small side network, achieving adaptation without backpropagating through the main network. BitFit (Ben Zaken et al., 2022) only tunes the bias of PLMs, which extremely reduces the trainable parameters. Diff pruning learns a sparse “diff” vector to modify a small percentage of pre-trained parameters, enabling efficient adaptation (Guo et al., 2021). Our proposed ADePT is an improved variant of PT, which can be also unified with other PEFT methods.
@@
Prompt tuning and its variants: PT (Lester et al., 2021) enables the adaptation of PLMs to downstream tasks by learning a soft prompt appended in front of model input. Prefix-tuning (Li & Liang, 2021) can be viewed as an extension of Prompt Tuning applied across the entire depth of the model. SPoT (Vu et al., 2022) utilizes trained prompts from source tasks as initialization, but it requires extensive search to find the optimal initialization. ATTEMPT (Asai et al., 2022) adopts an attention module to compose the trained prompts of source tasks. MPT (Wang et al., 2023) learns a transferable prompt by distilling from multiple task-specific prompts. XPrompt (Ma et al., 2022) empirically demonstrates the negative impact of trained prompt tokens and proposes a hierarchical structured pruning for a trained soft prompt, which can be seen as a post-training method. DPT (Xiao et al., 2023) adopts the product of two low-rank matrices to approximate the soft prompt. Residual Prompt Tuning (Razdaibiedina et al., 2023) passes the original soft prompt through a shallow network with a residual connection. Residual Prompt Tuning uses a shallow feed-forward neural network with a residual connection to reparameterize soft prompt embeddings, while ADePT uses a shallow feed-forward neural network to learn input token embedding offsets instead. DePT (Shi & Lipani, 2024) first discovers that the input token embedding offsets can enhance the performance of PT and uses short soft prompts to accelerate inference. ADePT can produce adaptive input token embedding offsets by shallow token-shared feed-forward neural networks, which can address the limitations of DePT as mentioned in Section 1. With a comparable number of trainable parameters, ADePT can outperform both vanilla PT and DePT.
@@ 3 Method
In this section, we first go back to the preliminaries of Prompt Tuning (PT) and Decomposed Prompt Tuning (DePT) in Section 3.1 and analyze the limitations of DePT in Section 3.2. Then, in Section 3.3, we give a detailed introduction of our proposed method, i.e., Adaptive Decomposed Prompt Tuning (ADePT). Finally, in Section 3.4, we give a theoretical analysis towards ADePT.
@@ 3.1 Preliminaries: Prompt Tuning (PT) and Decomposed Prompt Tuning (DePT)
Decomposed Prompt Tuning (DePT). DePT adapts PLMs to the new target downstream task via a short soft prompt $\bm{P}_{s_{1}}\in\mathbb{R}^{m_{s}\times d}$ and a pair of low-rank matrices, i.e., $\bm{A}\in R^{s\times r_{s}}$ and $\bm{B}\in\mathbb{R}^{r_{s}\times d}$, where $m$ denotes the length of short soft prompt and $r_{s}\ll\min(s,d)$ denotes the maximum rank of matrices $\bm{A}$ and $\bm{B}$. Similar to PT, the short soft prompt of DePT is prepended to the frozen input token embeddings. The product of low-rank matrices of DePT is used as the offsets of the input word embeddings. The loss of DePT for the target downstream task is formulated as:
@@
where the loss function $\mathcal{L}_{\mathrm{DePT}}$ is optimized only with respect to the short soft prompt $\bm{P_{s}}$ and the pair of low-rank matrices $\bm{A}$ and $\bm{B}$.
Table 1: The experimental results of DePT and ADePT on RTE and BoolQ tasks. All results are based on the T5-base model. The numbers within “DePT()” indicate the amount of cyclic left shift applied to the position-based
input embedding offsets of DePT. DePT (0) DePT (50) DePT(100) DePT(150) DePT(200) ADePT RTE 79.1 78.4 79.9 79.1 78.4 82.0 BoolQ 78.4 74.7 73.1 70.9 73.9 80.2
@@ 3.2 The Limitations of DePT
In this section, we explore several key factors that limit DePT performance.
@@
Table 2 reports the mean and variance of elements in input token embedding $\bm{e}$ and the elements in embedding offset $\Delta\bm{e}$ from DePT across two entire training datasets, i.e., RTE and BoolQ tasks. All results are calculated by their absolute values. We can observe that, the mean and variance of elements in $\Delta\bm{e}$ are only a few percent of those of elements in $\bm{e}$. For example, on the RTE task, the mean absolute value of elements in $\Delta\bm{e}$ is only 0.01, while the mean absolute value of elements in $\bm{e}$ is 6.07. This implies that DePT makes only minor changes to the input token embedding space, which may result in its inability to map the input token embeddings to the appropriate embedding space. This is because, for token $\bm{e}$, its offset can be any position $\Delta\bm{e}_{i}$, where $1\leq i\leq s$. The requirement that tokens within the same task should be unique causes the offsets of DePT to become extremely small, leading to the sub-optimization of DePT. The elements in the embedding offsets of ADePT have a much larger range of values than that of DePT. The optimal embedding space may lie outside the range of DePT, whereas ADePT may be able to access this embedding space.
@@ 3.3 Our method: Adaptive Decomposed Prompt Tuning (ADePT)
The limitations of DePT lie in its position-based token embedding offsets. Thus, to address this issue, the input token embedding offsets should be tailored for model input, and the corresponding embedding for each token should be unique after being offset. Surprisingly, we find that making the input token embedding offsets $\Delta\bm{E}$ a function of the input token embedding $\bm{E}$ can meet the above requirements. For an input token embedding $\bm{e}$, we want to get a function $f()$, which can produce the offset of this input token embedding, namely, $f(\bm{e})$. For the input token embeddings $\bm{E}=[\bm{e}_{1},\bm{e}_{2},\cdots,\bm{e}_{s}]$, the updated input token embedding can be formulated as $\bm{E}^{\prime}=\bm{E}+\Delta\bm{E}=[\bm{e}_{1}+f(\bm{e}_{1}),\bm{e}_{2}+f(\bm{e}_{s}),\cdots,\bm{e}_{s}+f(\bm{e}_{s})]$. For example, the input token embeddings $[\bm{a},\bm{b},\bm{c}]$ can be updated as $[\bm{a}+f(\bm{a}),\bm{b}+f(\bm{b}),\bm{c}+f(\bm{c})]$. Prepending the meaningless sequence $[\bm{t}_{1},\bm{t}_{2}]$ to input token embeddings $[\bm{a},\bm{b},\bm{c}]$, the updated input token embeddings are $[\bm{t}_{1}+f(\bm{t}_{1}),\bm{t}_{2}+f(\bm{t}_{2}),\bm{a}+f(\bm{a}),\bm{b}+f(\bm{b}),\bm{c}+f(\bm{c})]$. We can observe that the offsets for $[\bm{a},\bm{b},\bm{c}]$ are the same in this two scenario, i.e., $[f(\bm{a}),f(\bm{b}),f(\bm{c})]$. Therefore, if such a function $f$ exists, we can achieve input token embedding offsets that are tailored for model input, and the embedding for each token is unique within a task.
@@
To ensure that the number of trainable parameters does not exceed that of the vanilla PT, we use a short soft prompt $\bm{P_{s_{2}}}$, similar to DePT. The loss of ADePT is formulated as:
@@
where the loss function $\mathcal{L}_{\mathrm{ADePT}}$ is optimized only with respect to the short soft prompt $\bm{P_{s_{2}}}\in\mathbb{R}^{m\times d}$ and the parameters of the feed-forward neural network $\bm{W}_{\text{down}}$, $\bm{W}_{\text{up}}$, $\bm{b}_{1}$ and $\bm{b}_{2}$.
@@ 4.1 EXPERIMENTAL SETUP
Implementation Details. Following Shi & Lipani (2024), we use $100$ learnable virtual tokens as the soft prompt of PT. For our proposed ADePT, we adjust the hyperparameters to maintain an equivalent number of trainable parameters as PT. For instance, in the T5-base model, the token embedding dimension $d$ is $768$, so the number of trainable parameters is $l\times d=100\times 768=76800$. Following Shi & Lipani (2024), we search the length of soft prompt from $20$, $40$, $60$, and $80$. For ADePT, if using $60$ virtual tokens for soft prompt, the $d_{r}$ is got by solving the unequal equation $60\times 768+2\times d_{r}\times 768+d_{r}+768\leq 76800$. Thus, the $d_{r}\leq 19.49$ and $d$ is set to $19$ because the $d$ is the integer. According to this calculation method, the corresponding $d_{r}$ values for soft prompt lengths of $20$, $40$, $60$ and $80$ are $39$, $29$, $19$, and $9$, respectively. For a fair comparison of the T5-base model, we directly quote performance metrics from published papers (Mahabadi et al., 2021; Karimi Mahabadi et al., 2021; Asai et al., 2022; Wang et al., 2023; Sung et al., 2022; Shi & Lipani, 2024). For T5-3B model, we consistently use $60$ virtual tokens and bottleneck size $r=19$. Due to the lack of experimental results, for a fair comparison with the T5-3B model, we reproduce the experiments of the vanilla PT and DePT. For decoder-only PLMs, following Jain et al. (2024), we use $10$ virtual tokens for PT, $7$ virtual tokens and rank $r_{s}=3$ for DePT, $7$ virtual tokens and bottleneck size $r=1$ for ADePT, and rank $16$ for LoRA. For small datasets ($<70,000$ training samples) based on T5 model, we follow the learning strategy of Shi & Lipani (2024): we search the learning rate for the soft prompt from $13e-$ 1, $14e-$ 1, $15e-$ 1, and for the feed-forward neural network from $41e-$ 4, $51e-$ 5. For large datasets ($>70,000$ training samples) based on T5 model, we use learning rate 3e-1 for the soft prompt and $41e-$ 4 for the feed-forward neural networks. For the MBPP benchmark, following Jain et al. (2024), we use learning rates of $31e-$ 3 for the prompting-style tuning method, $41e-$ 4 for LoRA. For the few-shot learning, following the prior works (Asai et al., 2022; Wang et al., 2023; Shi & Lipani, 2024), we first pre-train five source tasks (i.e., MNLI, QQP, SST-2, SQUAD, and ReCoRD), and then select the best checkpoint to use as the initialization for few-shot training.

### v2

@@ Abstract
Prompt Tuning (PT) enables the adaptation of Pre-trained Large Language Models (PLMs) to downstream tasks by optimizing a small amount of soft virtual tokens, which are prepended to the input token embeddings. Recently, Decomposed Prompt Tuning (DePT) has demonstrated superior adaptation capabilities by decomposing the soft prompt into a shorter soft prompt and a pair of low-rank matrices. The product of the pair of low-rank matrices is added to the input token embeddings to offset them. Additionally, DePT achieves faster inference compared to PT due to the shorter soft prompt. However, in this paper, we find that the position-based token embedding offsets of DePT restrict its ability to generalize across diverse model inputs, and that the shared embedding offsets across many token embeddings result in sub-optimization. To tackle these issues, we introduce Adaptive Decomposed Prompt Tuning (ADePT), which is composed of a short soft prompt and a shallow token-shared feed-forward neural network. ADePT utilizes the token-shared feed-forward neural network to learn the embedding offsets for each token, enabling adaptive embedding offsets that vary according to the model input and better optimization of token embedding offsets. This enables ADePT to achieve superior adaptation performance without requiring more inference time or additional trainable parameters compared to vanilla PT and its variants. In comprehensive experiments across 23 natural language processing tasks and 4 typical PLMs of different scales, ADePT consistently surpasses the other leading parameter-efficient fine-tuning methods, and even outperforms the full fine-tuning in certain scenarios. We also provide a theoretical analysis towards ADePT. Code is available at https://github.com/HungerPWAY/ADePT.
@@ 1 Introduction
DePT also learns a short soft prompt $\bm{P}_{s}\in\mathbb{R}^{l_{s}\times d}$ with length $l_{s}$. The input token embeddings $\bm{E}$, after DePT processing, are formulated as $[\bm{P}_{s},\bm{E}+\bm{A}\bm{B}]$. In this paper, we focus on improving the fourth path by providing better offsets for the input token embeddings. Additionally, our improvement of the fourth path is orthogonal to the first three and can be integrated with their methods.
@@
  • Extensive evaluations across 23 NLP tasks and 4 PLMs of different scales show that ADePT surpasses leading PEFT methods, including full fine-tuning in certain scenarios. We theoretically analyze that ADePT possesses better expressive power than PT and DePT.
@@ 2 Related Works
Parameter-Efficient Fine-tuning: PEFT methods adapt PLMs to downstream tasks by optimizing a small number of parameters, significantly reducing computational resource demands. PEFT methods can be divided into four types:(1) Adapter and its variants (Houlsby et al., 2019; He et al., 2022; Rücklé et al., 2021; Ivison & Peters, 2022); (2) Low-rank Adaptation (LoRA) and its variants (Hu et al., 2021; Liu et al., 2022; Kopiczko et al., 2024); (3) Prompt Tuning (PT) and its variants (Lester et al., 2021; Li & Liang, 2021; Vu et al., 2022; Asai et al., 2022; Wang et al., 2023; Ma et al., 2022; Xiao et al., 2023; Razdaibiedina et al., 2023; Shi & Lipani, 2024); (4) other methods (Ben Zaken et al., 2022; Guo et al., 2021). The adapter inserts new modules into the transformer blocks (Houlsby et al., 2019). Hyperformer (He et al., 2022) uses a shared hypernetwork to generate task-conditioned adapters, reducing the trainable parameters in multi-task scenarios. AdapterDrop (Rücklé et al., 2021) removes adapters from lower transformer layers during training and inference, which reduces the computation overhead. Hyperdecoders (Ivison & Peters, 2022) generate input-specific adapters using a shared decoder for multi-task scenarios. LoRA (Houlsby et al., 2019) adopts the matrix product of a pair of low-rank matrices to approximate the updates of corresponding parameters. $(\text{IA})^{3}$ (Liu et al., 2022) rescales internal activations only using learned vectors injected into the attention and feedforward modules. Prompt tuning (PT) (Lester et al., 2021)adapts PLMs to the new downstream tasks by optimizing learnable virtual tokens. LST (Sung et al., 2022) reduces memory requirements by training a small side network, achieving adaptation without back-propagating through the main network. BitFit (Ben Zaken et al., 2022) only tunes the bias of PLMs, which extremely reduces the trainable parameters. Diff pruning learns a sparse “diff” vector to modify a small percentage of pre-trained parameters, enabling efficient adaptation (Guo et al., 2021). Our proposed ADePT is an improved variant of PT, which can also be unified with other PEFT methods.
@@
Prompt tuning and its variants: PT (Lester et al., 2021) enables the adaptation of PLMs to downstream tasks by learning a soft prompt appended in front of model input. Prefix-tuning (Li & Liang, 2021) can be viewed as an extension of Prompt Tuning applied across the entire depth of the model. SPoT (Vu et al., 2022) utilizes trained prompts from source tasks as initialization, but it requires extensive search to find the optimal initialization. ATTEMPT (Asai et al., 2022) adopts an attention module to compose the trained prompts of source tasks. MPT (Wang et al., 2023) learns a transferable prompt by distilling from multiple task-specific prompts. XPrompt (Ma et al., 2022) empirically demonstrates the negative impact of trained prompt tokens and proposes a hierarchical structured pruning for a trained soft prompt, which can be seen as a post-training method. DPT (Xiao et al., 2023) adopts the product of two low-rank matrices to approximate the soft prompt. Residual Prompt Tuning (Razdaibiedina et al., 2023) reparameterizes soft prompt embeddings with a shallow feed-forward network and residual connection, while ADePT uses the feed-forward network to learn input token embedding offsets—a distinct objective. DePT (Shi & Lipani, 2024) first discovers that the input token embedding offsets can enhance the performance of PT and uses short soft prompts to accelerate inference. ADePT produces adaptive input token embedding offsets via shallow token-shared networks, addressing the limitations of DePT mentioned in Section 1. With a comparable number of trainable parameters, ADePT can outperform both vanilla PT and DePT.
@@ 3 Method
In this section, we revisit the preliminaries of PT and DePT (Section 3.1) and analyze the limitations of DePT (Section 3.2). Then, we introduce our proposed Adaptive Decomposed Prompt Tuning (ADePT) in Section 3.3, followed by a theoretical analysis of ADePT in Section 3.4.
@@ 3.1 Preliminaries: Prompt Tuning (PT) and Decomposed Prompt Tuning (DePT)
Decomposed Prompt Tuning (DePT). DePT adapts PLMs to the new target downstream task via a short soft prompt $\bm{P}_{s_{1}}\in\mathbb{R}^{l_{s_{1}}\times d}$ and a pair of low-rank matrices, i.e., $\bm{A}\in R^{s\times r_{s}}$ and $\bm{B}\in\mathbb{R}^{r_{s}\times d}$, where $l_{s_{1}}$ denotes the length of short soft prompt and $r_{s}\ll\min(s,d)$ denotes the maximum rank of matrices $\bm{A}$ and $\bm{B}$. Similar to PT, the short soft prompt of DePT is prepended to the frozen input token embeddings. The product of low-rank matrices of DePT is used as the offsets of the input word embeddings. The loss of DePT for the target downstream task is formulated as:
@@
where the loss function $\mathcal{L}_{\mathrm{DePT}}$ is optimized only with respect to the short soft prompt $\bm{P}_{s_{1}}$ and the pair of low-rank matrices $\bm{A}$ and $\bm{B}$.
Table 1: The comparisons between DePT and ADePT on RTE and BoolQ tasks. All results are based on the T5-base model. The numbers within “DePT()” indicate the amount of cyclic left shift applied to the position-based input embedding offsets of DePT.
DePT (0) DePT (50) DePT(100) DePT(150) DePT(200) ADePT
RTE 79.1 78.4 79.9 79.1 78.4 82.0
BoolQ 78.4 74.7 73.1 70.9 73.9 80.2
@@ 3.2 The Limitations of DePT
In this section, we explore several key factors that limit the performance of DePT.
@@
Table 2 reports the mean and variance of elements in input token embedding $\bm{e}$ and the elements in embedding offset $\Delta\bm{e}$ from DePT across two entire training datasets, i.e., RTE and BoolQ tasks. All results are calculated by their absolute values. We can observe that the mean and variance of elements in $\Delta\bm{e}$ are only a few percent of those of elements in $\bm{e}$. For example, on the RTE task, the mean absolute value of elements in $\Delta\bm{e}$ is only 0.01, while the mean absolute value of elements in $\bm{e}$ is 6.07. This implies that DePT makes only minor changes to the input token embedding space, which may result in its inability to map the input token embeddings to the appropriate embedding space. This is because, for token $\bm{e}$, its offset can be any position $\Delta\bm{e}_{i}$, where $1\leq i\leq s$. The requirement that tokens within the same task should be unique causes the offsets of DePT to become extremely small, leading to the sub-optimization of DePT. The elements in the embedding offsets of ADePT have a much larger range of values than that of DePT. The optimal embedding space may lie outside the range of DePT, whereas ADePT may be able to access this embedding space.
@@ 3.3 Our method: Adaptive Decomposed Prompt Tuning (ADePT)
The limitations of DePT lie in its position-based token embedding offsets. To address this issue, the input token embedding offsets should be tailored for model input, and the corresponding embedding for each token should be unique after being offset. We find that making the input token embedding offsets $\Delta\bm{E}$ a function of the input token embedding $\bm{E}$ can meet such requirements. For an input token embedding $\bm{e}$, we want to get a function $f()$, which can produce the offset of this input token embedding, namely, $f(\bm{e})$. For the input token embeddings $\bm{E}=[\bm{e}_{1},\bm{e}_{2},\cdots,\bm{e}_{s}]$, the updated input token embedding can be formulated as $\bm{E}^{\prime}=\bm{E}+\Delta\bm{E}=[\bm{e}_{1}+f(\bm{e}_{1}),\bm{e}_{2}+f(\bm{e}_{s}),\cdots,\bm{e}_{s}+f(\bm{e}_{s})]$. For example, the input token embeddings $[\bm{a},\bm{b},\bm{c}]$ can be updated as $[\bm{a}+f(\bm{a}),\bm{b}+f(\bm{b}),\bm{c}+f(\bm{c})]$. Prepending the meaningless sequence $[\bm{t}_{1},\bm{t}_{2}]$ to input token embeddings $[\bm{a},\bm{b},\bm{c}]$, the updated input token embeddings are $[\bm{t}_{1}+f(\bm{t}_{1}),\bm{t}_{2}+f(\bm{t}_{2}),\bm{a}+f(\bm{a}),\bm{b}+f(\bm{b}),\bm{c}+f(\bm{c})]$. We can observe that the offsets for $[\bm{a},\bm{b},\bm{c}]$ are the same in this two scenario, i.e., $[f(\bm{a}),f(\bm{b}),f(\bm{c})]$. Therefore, if such a function $f$ exists, we can achieve input token embedding offsets that are tailored for model input, and the embedding for each token is unique within a task.
@@
To ensure that the number of trainable parameters does not exceed that of the vanilla PT, we use a short soft prompt $\bm{P}_{s_{2}}\in\mathbb{R}^{l_{s_{2}}\times d}$, similar to DePT. The loss of ADePT is formulated as:
@@
where the loss function $\mathcal{L}_{\mathrm{ADePT}}$ is optimized only with respect to the short soft prompt $\bm{P}_{s_{2}}$ and the parameters of the feed-forward neural network $\bm{W}_{\text{down}}$, $\bm{W}_{\text{up}}$, $\bm{b}_{1}$ and $\bm{b}_{2}$.
@@ 4.1 EXPERIMENTAL SETUP
Implementation Details. Following Shi & Lipani (2024), we use $100$ learnable virtual tokens as the soft prompt of PT. For our proposed ADePT, we adjust the hyperparameters to maintain an equivalent number of trainable parameters as PT. For instance, in the T5-base model, the token embedding dimension $d$ is $768$, so the number of trainable parameters is $l\times d=100\times 768=76800$. Following Shi & Lipani (2024), we search the length of soft prompt from $20$, $40$, $60$, and $80$. For ADePT, if using $60$ virtual tokens for soft prompt, the $d_{r}$ is got by solving the unequal equation $60\times 768+2\times d_{r}\times 768+d_{r}+768\leq 76800$. Thus, the $d_{r}\leq 19.49$ and $d$ is set to $19$ because the $d$ is the integer. According to this calculation method, the corresponding $d_{r}$ values for soft prompt lengths of $20$, $40$, $60$ and $80$ are $39$, $29$, $19$, and $9$, respectively. For a fair comparison of the T5-base model, we directly quote performance metrics from published papers (Mahabadi et al., 2021; Karimi Mahabadi et al., 2021; Asai et al., 2022; Wang et al., 2023; Sung et al., 2022; Shi & Lipani, 2024). For T5-3B model, we consistently use $60$ virtual tokens and bottleneck size $r=19$. Due to the lack of experimental results, for a fair comparison with the T5-3B model, we reproduce the experiments of the vanilla PT and DePT. For decoder-only PLMs, following Jain et al. (2024), we use $10$ virtual tokens for PT, $7$ virtual tokens and rank $r_{s}=3$ for DePT, $7$ virtual tokens and bottleneck size $r=1$ for ADePT, and rank $16$ for LoRA. For small datasets ($<70,000$ training samples) based on T5 model, we follow the learning strategy of Shi & Lipani (2024): we search the learning rate for the soft prompt from $13e-$ 1, $14e-$ 1, $15e-$ 1, and for the feed-forward neural network from $41e-$ 4, $51e-$ 5. For large datasets ($>70,000$ training samples) based on T5 model, we use learning rate 3e-1 for the soft prompt and $41e-$ 4 for the feed-forward neural networks. For the MBPP benchmark, following Jain et al. (2024), we use learning rates of $31e-$ 3 for the prompting-style tuning method, $41e-$ 4 for LoRA.